# Dealing with Uncertainty in Contextual Anomaly Detection

### v1

@@ 5 Conclusion
Through experiments on benchmark datasets and a real-world application in cardiology, we demonstrated that NS achieves state-of-the-art performance comparing with other contextual anomaly detection methods such as QCAD and ROCOD. Additionally, the ability to quantify uncertainty provides significant advantages in high-stakes domains like healthcare, where decision-making must often account for sparse or unreliable data.
Future work will explore extending the NS framework to other domains where contextual information plays a key role, like in the aortic diameters scenario. In addition, the case of vector-valued behavior needs to be addressed by taking into account the relationships among different behavioral variables. For example, the overall shape of the aorta might be anomalous even though every individual diameter (such as SoV or AA) may be normal. The use of multi-task gaussian processes (Bonilla, Chai, and Williams 2007) may be considered for this purpose.

### v2

@@ 5 Conclusion
Through experiments on benchmark datasets and a real-world application in cardiology, we demonstrated that NS achieves state-of-the-art performance comparing with other contextual anomaly detection methods such as QCAD and ROCOD. Additionally, the ability to quantify uncertainty provides significant advantages in high-stakes domains like healthcare, where decision-making must often account for sparse or unreliable data. Developing approaches that explicitly reason about predictive uncertainty is important for promoting trust in machine-learning systems. This is particularly true in high-stakes applications such as medicine, where predictions may influence decision-making and it is paramount to know when an output should be treated with caution or further investigated. Although NS sometimes offers only a modest margin over a plain Z‑score, our experiments are, to our knowledge, the first to reveal that such a simple baseline can already be remarkably competitive on contextual anomaly detection tasks, an observation that future benchmarks should not overlook.
Future work will explore extending the NS framework to other domains where contextual information plays a key role, like in the aortic diameters scenario. In addition, the case of vector-valued behavior needs to be addressed by taking into account the relationships among different behavioral variables. For example, the overall shape of the aorta might be anomalous even though every individual diameter (such as SoV or AA) may be normal. This can be addressed with multi-output/multi-task Gaussian processes (Bonilla et al., 2007), and more structured constructions that place Wishart process priors over context-dependent covariance matrices to explicitly capture correlations between behavioral variables (Heaukulani and van der Wilk, 2019).
@@
Acknowledgments
JD acknowledges the support of the Flemish government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme. A special thanks to Curzio Checcucci for his valuable feedback during the initial formulation of the problem.